What is the probability of at least one head?
If you throw a single fair coin multiple times, the probability of getting NO head is:For 1 throw: 1/2
For 2 throws: 1/2 squared = 1/4
For 3 throws: 1/2 cubed = 1/8
etc.
The probability of getting AT LEAST ONE head is the complement; for example, for 3 throws, it would be 1 minus 1/8.
What does probability measures?
From a scale of 1 to 0 probability measure an event will happen, maybe it will happen or it will not happen.
It is true.
What is The set of all possible outcomes in a probability experiment is the?
It is the outcome space.
In the Empirical Rule of data will fall in with two standard deviation.?
Approx 95% of the observations.
44%
If 4 are available, there are 24 possible ways to assign them. If 7 then there are 840.
What is the probability that the spinner and number cube both land on 2?
That would depend on how many numbers are on the spinner and the cube. The more numbers there are, the less likely it is that they would both land an any given number.
What is the probability of observing a sample mean of 18 or more from a sample size of 35?
Without knowing the data which is being sampled, it is impossible to answer this other than by saying that the probability is between 0 and 1 inclusive.
Consider a company. If you sample the annual pay of the employees, any mean will be greater than 18 as everyone will be taking home more than £18 per year, so the probability is 1.
Consider a school. If you sample the lengths of feet of the pupils, any mean will be less than 18 as all the feet are less than 18 inches long, so the probability is 0.
If an event has a probability of 1 then it is?
Then it is a certainty that the event is 100% to happen.
4/13
As the two events are independent multiply the respective probabilities together:
pr(5) = 1/6
pr(tails) = 1/2*
→ pr(5 and tails) = 1/6 × 1/2 = 1/12
*Actually it is slightly less than 1/2 as there is a very small probability that the coin will end up on its edge; for a small coin it is so small that it can be ignored, but with a scaled up coin it becomes a real possibility as (accidentally) demonstrated in the 1997 Royal Institution Christmas Lecture "The Magical Maze"
The possible prime numbers are 2, 3, 5
The possible numbers greater than 3 are 4, 5, 6
→ a success is 2 3, 4, 5, 6
→ pr(prime or > 3) = 5/6
one in three
1 2 = under 3
3 4 5 6 = at least 3
so 4/6 probability
coin is 1 in 2
4/6 x 1/2 = 4/12 = 1/3
Pr(Both picked 3) = 1/16Pr(Both picked the same) = 1/4.
Assuming you want the probability FOR A SINGLE TRY, and you want the numbers in that exact order, the probability for each part (for instance, first = red; or second = green) is 1/4; therefore, the probability for the combination is (1/4) to the power 4.
What the answer
If you throw the fair die once then the chances are two as there are two numbers to be shown that is 3 or 5.
Total number of faces and numbers in dice = 6.
So, Probability of two numbers falling = 2/6 = 1/3.
If you roll the same die 180 times then chances of landing on either 3 or 5 = 180 x 1/3 = 60.
You need to ask 1 question at a time, rewording it into 200 characters or less - it helps you to actually think about what is required instead of getting others to do your work for you.
z = (x - µ)/standard deviation
= (52 - 50)/20
= 0.1
(If z is negative it is below the mean.)
This is supposed to be Y > u
What is a example of cluster sampling?
Cluster sampling is a scheme which is used when sampling from the whole population would be too expensive - in terms of time or money. The sample space is divided into clusters, a random sample of clusters is selected and then every member of the population within the selected clusters is studied.
For example, suppose you wanted to collect information from schools across a country and had calculated that a 5% sample was required. Rather than criss-crossing the country, you could divide the country into units: for example counties. You then select counties (your clusters) so that they cover 5% of the nation's schools. Visit each chosen county and sample all schools in it. The selection of counties would probably also need to be controlled so that urban and rural areas are properly represented.